# Probability of Differentiation Reveals Brittleness of Homogeneity Bias in Large Language Models

### v1

@@ Abstract
Homogeneity bias in Large Language Models (LLMs) refers to their tendency to homogenize the representations of some groups compared to others. Previous studies documenting this bias have predominantly used encoder models, which may have inadvertently introduced biases. To address this limitation, we prompted GPT-4 to generate single word/expression completions associated with 18 situation cues – specific, measurable elements of environments that influence how individuals perceive situations and compared the variability of these completions using probability of differentiation. This approach directly assessed homogeneity bias from the model’s outputs, bypassing encoder models. Across five studies, we find that homogeneity bias is highly volatile across situation cues and writing prompts, suggesting that the bias observed in past work may reflect those within encoder models rather than LLMs. Furthermore, these results suggest that homogeneity bias in LLMs is brittle, as even minor and arbitrary changes in prompts can significantly alter the expression of biases. Future work should further explore how variations in syntactic features and topic choices in longer text generations influence homogeneity bias in LLMs.
@@ 1 Introduction
Recent research has increasingly focused on the tendency of LLMs to homogenize the representations of minority groups. Studies such as Cheng et al. (2023a) and Cheng et al. (2023b) observed that LLMs portray minority groups with positive yet homogeneous narratives. Building on these findings, Lee et al. (2024) introduced the concept of homogeneity bias – a sophisticated bias that impacts not only the narrative content but also the manner in which groups are discussed. They found that ChatGPT portrayed subordinate racial/ethnic and gender groups as more homogeneous than their dominant group counterparts, potentially promoting erasure and stereotyping.
@@
To address this limitation, we propose a novel and complementary method to assess homogeneity bias in LLMs that does not rely on encoder models. Our approach involves two steps: First, we use single word or expression completion prompts focusing on various human activities. For example, we ask the model to complete a sentence about a sport that an African American man is playing. Second, we quantify the variability of these completions using probability of differentiation, a measure commonly used in social psychology to quantify how humans perceive variability of groups (Linville et al., 1989; Park and Judd, 1990; Simon and Pettigrew, 1990; Judd et al., 1991). Probability of differentiation calculates the likelihood that two randomly chosen completions for a writing prompt will differ, with a higher value indicating greater heterogeneity. By using this approach, we can directly assess homogeneity bias from the model’s outputs.
@@ 1.1 The Present Research
Using this method, we compared the variability of human activities associated with eight groups at the intersection of four racial/ethnic and two gender groups across 18 different human activities. We expected that probability of differentiation of socially subordinate groups would be consistently smaller than that of dominant groups across all 18 human activities. However, we found that this is not the case. Rather, our initial study found homogeneity bias with regards to both race/ethnicity and gender varied greatly depending on the topic of study. Homogeneity bias remained variable in subsequent ablation studies using a different model version and an alternative identity signaling method. Our findings challenge the assumption that LLMs represent subordinate groups with greater homogeneity across every measure of homogeneity and suggest that homogeneity bias many manifest in other forms, such as through syntactic elements, which are not captured by single word or expression completions. This calls for a need for future research to explore other ways in which homogeneity bias manifest in LLMs.
@@ 2.1 Name Selection
In our writing prompts, we used names to signal group identities, representing eight groups at the intersection of four racial/ethnic (i.e., African, Asian, Hispanic, and White American) and two gender (i.e., men and women) group identities. Utilizing the Name-Trait Perceptions dataset (Elder and Hayes, 2023), which comprises of 1,000 common American first names rated with respect to group identities (i.e., race/ethnicity and gender) and traits (e.g., aggressive, hardworking), we randomly sampled 15 first names for each intersectional group. [^1] Notably, the dataset had limited Asian names, with 37 names identified as Asian, 16 of which were of Asian (American) men.
@@ 2.2 Completion Prompts
For each of the 18 cues, we designed instructions and writing prompts. The instructions were supplied as a system message which helped determine the behavior of GPT-4. By default, the system message contained, “Complete the following sentence with a single word or expression. Only return the word or expression." and was followed by the instructions in Table A2. The writing prompt was supplied as a user message. Then, using the OpenAI API, we had GPT-4 (gpt-4-0125-preview; 16 February 2024) complete these prompts. We generated fifty completions for each name, totaling 6,000 completions for each cue.
Occasionally, GPT-4 indicated that there wasn’t sufficient information to respond or generated text unrelated to the prompt (i.e., non-compliances). These were often longer than typical responses. To identify these, we counted the number of words in each response, manually inspected the lengthier responses, and removed noncompliances. The number of non-compliances are reported in Table A16 of the Supplementary Materials.
2.3 Probability of Differentiation
To assess the variability of completions for each group within a situation cue, we computed probability of differentiation. This measure, used in the perceived variability literature to evaluate phenomena like the out-group homogeneity effect (Linville et al., 1989; Park and Judd, 1990; Simon and Pettigrew, 1990; Judd et al., 1991), quantifies the likelihood that two randomly selected responses will be different from each other (see Equation 1).
$P_{d}=1-\sum_{i=1}^{m}p_{i}^{2}$ (1)
In the Equation, $p_{i}$ denotes the proportion of completions corresponding to the $i$th response category, and $m$ represents the total number of unique response categories. This metric is appropriate for assessing homogeneity in LLM-generated text for the following reasons: (1) The metric quantifies the variation in non-numeric, categorical variables like jobs or sports, allowing for the quantification of variability associated with different groups; (2) The measure increases when completions are more evenly distributed across categories, leading to lower values for groups frequently linked to a predominant category – stereotyping – and higher values for groups without a dominant association. For a more detailed discussion of the measure, see Linville et al. (1989).
2.4 Preprocessing
2.5 Cluster Bootstrapping
We implemented cluster bootstrapping to reliably compare probability of differentiation values across groups and to assess the uncertainty of the measure. This method is well-suited for datasets where individual observations are organized into clusters (Huang, 2018). In our dataset, completions associated with each racial/ethnic and gender group were nested within names. Cluster bootstrapping helped estimate the variability and confidence intervals of our metric by accounting for the clustered structure of our data. Unlike regular bootstrapping where individual observations within a group are resampled, cluster bootstrapping resamples the cluster themselves. Specifically, we resampled the names within each racial/ethnic or gender group, including all observations linked to each resampled name to compute $P_{d}$. We repeated this 1,000 times to establish 95% confidence intervals (95% CIs).
2.6 Meta-Analysis
To assess the consistency of homogeneity bias across situation cues for each group comparison, we first calculated Cohen’s d effect sizes for each group comparison within individual situation cues. We then conducted a random-effects meta-analysis using the meta package in R (R Version 4.4.0; $K$ = 18). We chose random-effects models because we expected the effect of race and gender to differ across situation cues. In addition to reporting the meta-analytic estimates for each group comparison, we conducted tests of heterogeneity and reported $I^{2}$ statistics. These tests demonstrated that a random-effects model was more appropriate for our analysis and allowed us to quantify the variability of effect sizes across situation cues. We refer to this study as the Main Study.
3 Results
3.1 Race
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and African Americans ($I^{2}$ = 99.92%, p <.$<.001$). Compared to African Americans, White Americans were presented more homogeneously in 8 situations, less homogeneously in 10 situations, and similarly in 0 situations. Effect sizes were as high as $d=2.54$ (African Americans more homogeneous) in Grooming, $t(1266.94)=56.84$, p <.$<.001$, and as low as $d=-6.10$ (White Americans more homogeneous) in Commuting, $t(1592.92)=-136.41$, p <.$<.001$. Collapsing across this heterogeneity, there was no overall difference between White and African Americans in the probability of differentiation, ($d=-0.86$, 95% CI = [$-2.02,0.30$]).
There was also incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and Asian Americans ($I^{2}$ = 99.92%, p <.$<.001$). Compared to Asian Americans, White Americans were presented more homogeneously in 10 situations, less homogeneously in 0 situations, and similarly in 8 situations. Effect sizes were as high as $d=5.10$ (Asian Americans more homogeneous) in Telephone, $t(1527.62)=114.12$, p <.$<.001$, and as low as $d=-3.82$ (White Americans more homogeneous) in Exam, $t(1318.79)=-85.44$, p <.$<.001$. Collapsing across this heterogeneity, there was no overall difference between White and Asian Americans in the probability of differentiation, ($d=0.29$, 95% CI = [$-0.78,1.35$]).
Consistent with other contrasts, we also observed incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and Hispanic Americans ($I^{2}$ = 99.88%, p <.$<.001$). Compared to Hispanic Americans, White Americans were presented more homogeneously in 7 situations, less homogeneously in 9 situations, and similarly in 2 situations. Effect sizes were as high as $d=3.91$ (Hispanic Americans more homogeneous) in Telephone, $t(1792.35)=87.39$, p <.$<.001$, and as low as $d=-3.67$ (White Americans more homogeneous) in Commuting, $t(1592.65)=-82.04$, p <.$<.001$. Collapsing across this heterogeneity, there was no overall difference between White and Hispanic Americans in the probability of differentiation, ($d=-0.47$, 95% CI = [$-1.35,0.41$]). See Table A4 and Figure 1.
3.2 Gender
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of men and women ($I^{2}$ = 99.94%, p <.$<.001$). Compared to women, men were presented more homogeneously in 9 situations, less homogeneously in 8 situations, and similarly in 1 situations. Effect sizes were as high as $d=11.26$ (women more homogeneous) in Online, $t(1934.78)=251.67$, p <.$<.001$, and as low as $d=-6.23$ (men more homogeneous) in Commuting, $t(1597.14)=-139.40$, p <.$<.001$. Collapsing across this heterogeneity, there was no overall difference between men and women in the probability of differentiation, ($d=0.73$, 95% CI = [$-1.25,2.70$]).
4 Ablation Studies
Previous work in the literature had documented evidence of homogeneity bias in LLMs. However, our study did not corroborate the presence of such bias; there was incredibly high heterogeneity in effect sizes comparing the probability of differentiation between dominant and subordinate groups. We propose explanations for this discrepancy, which could be due to variations in the model version (i.e., GPT-4 versus GPT-3.5), the method used to signal group identity (i.e., names versus group labels), and the specificity of the prompts (i.e., specific versus general prompts). To explore whether these factors might explain the lack of consistent homogeneity bias in our findings, we first conducted three ablation studies. These studies assess each variable – model version, identity signaling method, and prompt specificity – separately to determine their individual contributions to the observed variations in bias. This approach helps pinpoint the underlying reasons for the difference in findings and clarifies the conditions under which the bias manifests.
For the first two ablation studies where we used the same situation cues, we examined if each specific effect in the Main study replicated in the ablation studies. Following the practices of the Reproducibility Project Open Science Collaboration (2015), we first transformed the $t$ statistics for each comparison into correlation coefficients, calculated the proportion of study-pairs where the effect of the Main study was in the CI of the ablation study effect, then compared this with the expected proportion that the ablation studies would replicate using a goodness-of-fit $\chi^{2}$ test.
The effects of the ablation studies did not replicate those of the Main Study. To test if variations in homogeneity bias are a general feature of LLMs, we conducted a fourth and final ablation study, assessing replicability after making minimal modification to the prompts.
4.1 GPT-4 or GPT-3.5
Previous research utilized gpt-3.5-turbo for data collection, whereas our study implemented gpt-4-0125-preview. Newer models like GPT-4 often incorporate enhanced safety features and mitigation strategies to reduce bias, following advancements in algorithmic fairness and more diverse training data. To examine if these improvements contributed to diminished homogeneity bias, we conducted an ablation study using gpt-3.5-turbo. Finding evidence of bias in the ablation study would indicate that improvements in GPT-4 may explain the variations in our findings. We refer to this study as the GPT-3.5 Study.
Figure 3: Probability of Differentiation of racial/ethnic groups for the GPT-3.5 Study.
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and African, Asian, and Hispanic Americans ($I^{2}$s = 99.93%, 99.95%, and 99.91%, ps <.$<.001$). Collapsing across this heterogeneity, there was no overall difference in the probability of differentiation of White and African Americans ($d=-0.37$, 95% CI = [$-1.87,1.13$]), White and Asian Americans ($d=0.24$, 95% CI = [$-1.85,2.33$]), and White and Hispanic Americans ($d=-0.40$, 95% CI = [$-1.51,0.71$]). See Table A9 and Figure 3. Similarly, there was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of men and women ($I^{2}$ = 99.95%, p <.$<.001$). Collapsing across this heterogeneity, there was no overall difference between men and women in the probability of differentiation, ($d=-0.55$, 95% CI = [$-3.59,2.49$]). See Table A10 and Figure 4. Furthermore, the effects of the ablation study did not replicate those of the Main Study. Of the 72 group comparisons, only one (1.39%) of the GPT-3.5 Study CIs contained the Main Study effect size (significantly lower than the expected value of 83.4%, p <.$<.001$).
Figure 4: Probability of Differentiation of gender groups for the GPT-3.5 Study.
4.2 Names or Group Labels
Previous research signaled group identity using single group labels (e.g., Hispanic American men), while our approach involved using collections of names. Names, as distinct and personal identifiers, could evoke more detailed and varied representations of individuals within groups, potentially reducing stereotypical portrayals. On the other hand, single group labels may promote more generic and homogenized representations, focusing on collecting characteristics rather than individual diversity. This focus may increase the model’s reliance on stereotypical traits, thereby enhancing homogeneity bias. To investigate if using names attenuates homogeneity bias, we conducted an ablation study using group labels to signal group identity. As completions associated with each racial/ethnic and gender group were no longer nested within names, we performed regular bootstrapping to derive 95% CIs. Conducting this comparison helped determine if the method of signaling group identity influenced homogeneity bias. Finding evidence of bias in the ablation study would indicate that variations in group identity signaling methods may explain the variations in our findings. We refer to this study as the Group Labels Study.
Figure 5: Probability of Differentiation of racial/ethnic groups for the Group Labels Study.
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and African, Asian, and Hispanic Americans ($I^{2}$s = 99.96%, 99.96%, and 99.96%, ps <.$<.001$). Collapsing across this heterogeneity, there was no overall difference in the probability of differentiation of White and African Americans ($d=0.33$, 95% CI = [$-3.66,4.32$]), White and Asian Americans ($d=5.27$, 95% CI = [$-0.14,10.68$]), and White and Hispanic Americans ($d=3.95$, 95% CI = [$-1.65,9.55$]). See Table A7 and Figure 5. Similarly, there was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of men and women ($I^{2}$ = 99.95%, p <.$<.001$). Collapsing across this heterogeneity, there was no overall difference between men and women in the probability of differentiation, ($d=1.29$, 95% CI = [$-1.85,4.44$]). See Table A8 and Figure 6. Furthermore, the effects of the ablation study did not replicate those of the Main Study. Of the 72 group comparisons, only one (1.39%) of the Group Labels Study CIs contained the Main Study effect size (significantly lower than the expected value of 83.4%, p <.$<.001$).
Figure 6: Probability of Differentiation of gender groups for the Group Labels Study.
4.3 Prompt Specificity
Another potential factor contributing to the absence of homogeneity bias is prompt specificity. Prior research by Cheng et al. (2023a) found that LLMs tend to amplify stereotypical characteristics of groups in response to more general writing prompts. To investigate if specificity of prompts affects the manifestation of bias, we designed writing prompts that were more general relative to the ones focusing on individual situation cues. This approach allowed us to assess whether more general prompts led to more pronounced bias, thereby providing insights into how prompt specificity impacts homogeneity bias.
We designed writing prompts that would grant the LLM the flexibility to generate responses that weren’t constrained to a single situation cue. These writing prompts can be found in Table A11. The expectation was that when the model is given more flexibility, it would exhibit more homogeneity for subordinate groups. The instructions were supplied as a system message which helped determine the behavior of GPT-4. By default, the system message contained, “Complete the following sentence with a single word or expression. Only return the word or expression." The writing prompt was supplied as a user message. Then, using the OpenAI API, we had GPT-4 (gpt-4-0125-preview; 2 May 2024) complete these prompts. We generated fifty completions for each name, totaling 6,000 completions for each writing prompt. We refer to this study as the General Prompts Study.
Figure 7: Probability of Differentiation of racial/ethnic groups for the General Prompts Study.
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and African, Asian, and Hispanic Americans ($I^{2}$s = 99.86%, 99.96%, and 99.91%, ps <.$<.001$). Collapsing across this heterogeneity, there was no overall difference in the probability of differentiation of White and African Americans ($d=0.56$, 95% CI = [$-0.28,1.39$]), White and Asian Americans ($d=-0.74$, 95% CI = [$-3.08,1.60$]), and White and Hispanic Americans ($d=0.07$, 95% CI = [$-0.90,1.03$]). See Table A12 and Figure 7). Similarly, there was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of men and women ($I^{2}$ = 99.78%, p <.$<.001$). Collapsing across this heterogeneity, men had significantly higher probability of differentiation than women ($d=1.52$, 95% CI = [$0.99,2.05$]). See Table A13 and Figure 8).
Figure 8: Probability of Differentiation of gender groups for the General Prompts Study.
4.4 Minimal Prompt Modification
The effects of the ablation studies did not replicate those of the Main Study and raised the possibility that variation in homogeneity bias is a general feature of LLMs rather than a feature specific to the groups we studied. To test this possibility, we made a minimal modification to the writing prompts in Table A2 replacing the word “person" with “individual." We then assessed replicability using a goodness-of-fit $\chi^{2}$ test. If homogeneity bias did not vary much in this study, that would suggest Main Study effects are specific to the groups studied. On the other hand, if homogeneity bias varied greatly like in the Main Study, that would suggest that the phenomenon is a broader feature of LLMs. We refer to this study as the Individual Prompt Study.
Figure 9: Probability of Differentiation of racial/ethnic groups for the Individual Prompt Study.
There was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of White and African, Asian, and Hispanic Americans ($I^{2}$s = 99.93%, 99.92%, and 99.88%, ps <.$<.001$). Collapsing across this heterogeneity, there was no overall difference in the probability of differentiation of White and African Americans ($d=-0.41$, 95% CI = [$-1.60,0.77$]), White and Asian Americans ($d=0.76$, 95% CI = [$-0.28,1.80$]), and White and Hispanic Americans ($d=-0.22$, 95% CI = [$-1.08,0.64$]). See Table A14 and Figure 9). Similarly, there was incredibly high heterogeneity in effect sizes comparing the probability of differentiation of men and women ($I^{2}$ = 99.96%, p <.$<.001$). Collapsing across this heterogeneity, there was no overall difference between men and women in the probability of differentiation, ($d=-1.15$, 95% CI = [$-3.00,0.70$]). See Table A15 and Figure 10. Furthermore, the effects of the ablation study did not replicate those of the Main Study. Of the 72 group comparisons, seven (9.72%) of the Individual Prompt Study CIs contained the Main Study effect size (significantly lower than the expected value of 83.4%, p <.$<.001$).
Figure 10: Probability of Differentiation of gender groups for the Individual Prompt Study.
5 Discussion
5.1 Homogeneity Bias is Brittle
We found that the dominant racial/ethnic and gender groups were not consistently associated with more diverse human experiences than their subordinate group counterparts. Instead, relative heterogeneity varied significantly across situation cues, which were underscored by the consistently high $I^{2}$ statistics. Furthermore, the findings in the Main Study did not replicate across subsequent ablation studies where homogeneity bias remained highly variable but the consistency of group differences varied with the prompt. These findings align with previous observations that the behavior of LLMs is highly sensitive to the prompt used (e.g., Lu et al., 2022; Srivastava et al., 2023; Sclar et al., 2023; Pezeshkpour and Hruschka, 2023). Our results indicate that homogeneity bias in LLMs, as measured by probability of differentiation, is brittle, with minor and arbitrary changes in prompts significantly altering outcomes.
5.2 Limitations and Future Work
Despite extensive efforts to control for other confounding factors, not all potential confounds were accounted for. Homogeneity bias may manifest in longer forms of text generation, such as storytelling, but not in single word/expression completions. This could be because, although LLMs may associate two groups with equally diverse human activities, they might exhibit greater variations in narrative style, such as in diction and syntactic structures. These variations might be captured using sentence embeddings but not through probability of differentiation on single word/phrase completions. Future work should explore homogeneity bias in terms of syntactic features to better understand their role in the bias.
Moreover, homogeneity bias may be more apparent in longer text generations because models might select topics where homogeneity bias is more profound. It is possible that under less constrained conditions – such as in response to more general writing prompts – LLMs may naturally gravitate towards topics where the bias is more pronounced. For instance, in our Main study and subsequent ablation studies, White Americans were consistently associated with a wider variety of sports compared to other racial/ethnic groups. If LLMs disproportionately rely on topics like sports for text generation, this would not only explain homogeneity bias documented in past research but also explain evidence of homogeneity found in the General Prompts study with regards to gender. Future work should investigate how the choice of topics and the breadth of content allowed in prompts influence the manifestation of homogeneity bias in LLMs.
6 Conclusion
Our study proposed a novel method to assessing homogeneity bias in LLMs by utilizing a single word/expression completion prompts and a direct measure of variability derived from social psychology. This method allows us to assess biases directly from the LLM outputs, without the confounding influence of encoder models. Our findings reveal that homogeneity bias is highly volatile across situation cues and writing prompts, with most effects not replicating in an ablation study where only one word in the writing prompt was replaced. These results suggest that homogeneity bias observed in past work may actually have stemmed from the encoder models they used and that homogeneity bias may be brittle. Finally, we propose that stylistic elements, such as diction and syntactic structure, as well as the models’ reliance on specific topics, like sports, where bias appears more pronounced in favor of dominant groups, could influence homogeneity bias. To ensure fair and equitable representations of groups in LLM-generated texts, future research must explore if and how the proposed elements contribute to the bias.
Appendix A Supplementary Materials
A.1 Pilot Study
In the pilot study, we came up with eight areas of human experience and designed instructions and writing prompts for each (see Table A3 for the instructions and prompts). The instructions were supplied as a system message, and the writing prompt was supplied as a user message. By default, the system message was, “You are a chatbot. Don’t generate notes. If you cannot determine the answer, guess," followed by the instructions outlined in the Table. We then employed the OpenAI API, specifically GPT-4 (gpt-4-0125-preview), to complete these prompts. We generated 50 completions for each name, resulting in a total of 6,000 completions for each human activity. The numbers of non-compliances in the pilot study are reported in Table A1.
Non-compliances
Car 0
Festival 61
Food 0
Hobby 0
Job 14
Major 6
Music 2
State 61
Total 144
A.1.1 Results
A.1.2 Conclusion
Table A2: The 18 situation cues within the objects, events, and activities domain, along with their corresponding instructions and writing prompts.

### v2

# Probability of Differentiation Reveals Brittleness of Homogeneity Bias in GPT-4

@@ Abstract
Homogeneity bias in Large Language Models (LLMs) refers to their tendency to homogenize the representations of some groups compared to others. Previous studies documenting this bias have predominantly used encoder models, which may have inadvertently introduced biases. To address this limitation, we prompted GPT-4 to generate single word/expression completions associated with 18 situation cues—specific, measurable elements of environments that influence how individuals perceive situations and compared the variability of these completions using probability of differentiation. This approach directly assessed homogeneity bias from the model’s outputs, bypassing encoder models. Across five studies, we find that homogeneity bias is highly volatile across situation cues and writing prompts, suggesting that the bias observed in past work may reflect those within encoder models rather than LLMs. Furthermore, we find that homogeneity bias in LLMs is brittle, as even minor and arbitrary changes in prompts can significantly alter the expression of biases. Future work should further explore how variations in syntactic features and topic choices in longer text generations influence homogeneity bias in LLMs.
Probability of Differentiation Reveals Brittleness of Homogeneity Bias in GPT-4
Messi H.J. Lee Washington University in St. Louis St. Louis, Missouri 63130 hojunlee@wustl.edu                        Calvin K. Lai Rutgers University New Brunswick, NJ 08901 calvin.lai@rutgers.edu
@@ 1 Introduction
1.1 Homogeneity Bias in AI
Building on these concerns, recent research on bias in LLMs has begun to focus on homogeneity bias—a form of stereotyping where AI models represent certain groups as more uniform than others. For example, Lee et al. (2024) found that texts generated by a state-of-the-art LLM about racial minorities in the U.S. and women were more homogeneous than those about White Americans and men. Cheng et al. (2023b) reported similar findings, highlighting that LLM simulations of marginalized groups are more susceptible to caricature—an exaggerated narrative of the demographic group.
Homogeneity bias in AI models, including LLMs, has significant implications for social representation and equity. As AI increasingly serves as a key source of social information, groups affected by this bias face risks of cultural erasure-the marginalization and suppression of their identities and histories-often driven by under-representation and stereotypical portrayals of marginalized communities (e.g., Kelly, 2017). Homogeneous, stereotype-based representations distort understanding of group identities and experiences, with research showing that repeated exposure to such portrayals shapes attitudes toward social groups (e.g., Park et al., 2007; Saleem et al., 2017). Moreover, evidence suggests that biases in AI can influence political decision-making (Fisher et al., 2024), raising concerns that homogeneity bias in AI outputs could perpetuate stereotypes and discrimination, amplifying systemic inequities.
1.2 Past Methods to Assess Homogeneity Bias
@@
Figure 1: A visualization of the study design. A completion prompt is supplied to the LLM, and the completions are used to compute probability of differentiation. In this example, completions for “Victoria" (top) are more evenly distributed across sports categories, yielding a higher probability of differentiation (0.80), while those for “Tanisha" (bottom) are more concentrated, resulting in a lower value (0.55).
@@ 1.3 The Present Research
To address this limitation, we propose a complementary method to assess homogeneity bias in LLMs that does not rely on encoder models (see Figure 1). Our approach involves two steps: First, we use single word or expression completion prompts focusing on various human activities. For example, we ask the model to complete a sentence about a sport that an African American man is playing. Second, we quantify the variability of these completions using probability of differentiation, a measure commonly used in social psychology to quantify how humans perceive variability of groups (Linville et al., 1989; Park and Judd, 1990; Simon and Pettigrew, 1990; Judd et al., 1991). Probability of differentiation calculates the likelihood that two randomly chosen completions for a writing prompt will differ, with a higher value indicating greater heterogeneity. By using this approach, we directly assess homogeneity bias from the model’s outputs.
Using this method, we compared the variability of human activities associated with eight groups at the intersection of four racial/ethnic and two gender groups across 18 different human activities. We expected that probability of differentiation of socially subordinate groups would be consistently smaller than that of dominant groups across all 18 human activities. However, we found that this is not the case. Rather, we found that homogeneity bias varied greatly depending on the topic of study. Homogeneity bias remained variable in subsequent ablation studies using a different model version and an alternative identity signaling method. Our findings challenge the assumption that LLMs associate subordinate groups with greater homogeneity across every measure of homogeneity and suggest that homogeneity bias may manifest in subtle forms, such as through syntactic elements, which cannot be captured from single word or expression completions. This calls for a need for future research exploring ways in which homogeneity bias manifests in LLMs.
@@ 2 Experiment
This section outlines the experimental design. All aspects of the experiment described here were consistent across all studies conducted.
@@ 2.1 Name Selection
In our writing prompts, we used names to signal group identities, representing eight groups at the intersection of four racial/ethnic (i.e., African, Asian, Hispanic, and White American) and two gender (i.e., men and women) groups. Utilizing the Name-Trait Perceptions dataset (Elder and Hayes, 2023), which comprises of 1,000 common American first names rated with respect to group identities (i.e., race/ethnicity and gender) and traits (e.g., hardworking). We randomly sampled 15 names per group to ensure a robust representation of intersectional identities. While coverage for smaller groups is inherently limited, [^1] For instance, only 37 names were identified as Asian, 16 belonging to Asian (American) men. this dataset is an effective tool to signal group identities in alignment with U.S. demographics, offering both breadth and reliability in intersectional representation.
@@ 2.2 Completion Prompts
For each of the 18 cues, we designed instructions and writing prompts. The instructions were supplied as a system message which helped determine the behavior of GPT-4. By default, the system message contained, “Complete the following sentence with a single word or expression. Only return the word or expression." and was followed by the instructions in Table 1. The writing prompt was supplied as a user message. Then, using the OpenAI API, we had GPT-4 (gpt-4-0125-preview; 16 February 2024) complete these prompts. We generated 50 completions for each name, totaling 6,000 completions for each cue.
Past methods, such as Lee et al. (2024), used contextualized embeddings to measure homogeneity bias. However, this approach made it difficult to disentangle whether the similarity arose from textual elements like syntax and word choice or from patterns in human experiences associated with different groups. The controlled design of this work addresses that limitation by focusing on single-word or expression completions. By removing confounding textual elements, it enables a more precise examination of the homogeneity tied specifically to group experiences.
Table 1: The 18 situation cues within the objects, events, and activities domain (Rauthmann et al., 2014; Saucier et al., 2007; Pervin, 1978), along with their corresponding instructions and writing prompts.